-Predict where the largest jump between successive ionization energies occurs for Li.

-between the second and third ionization energies

-between the fourth and fifth ionization energies

-between the first and second ionization energies

-between the third and fourth ionization energies

-Predict where the largest jump between successive ionization energies occurs for Be.

-between the second and third ionization energies

-between the first and second ionization energies

-between the third and fourth ionization energies

-between the fourth and fifth ionization energies
Answer #1

Here is the solution of your question. If you have any doubt or need any clarification please comment in comment box and will definitely resolve your query. If you find useful please upvote it. Thanks in advance.

ha I. E = -13.6x3 = -122.4 22 N -4 NS formula used I.E = -13-6x2 evlatom for Li et z =3 | 4.= Ez-Ez = -13.6-(-30-6)= 17ev y =
Answer #2

Li= first and second

Be= second and third

answered by: yasna